# FILP-3D: Enhancing 3D Few-shot Class-incremental Learning with Pre-trained Vision-Language Models

### v1

@@ 4 Benckmark for 3D FSCIL task
Studies on 3D FSCIL benchmarks are still in early stage. The sole benchmark introduced by Chowdhury et al. [3] has several notable limitations: 1) In the synthetic data to synthetic data (S2S) task set forth by Chowdhury et al. [3], the total number of classes is limited. This restricts the capacity of the model for incremental learning and makes it challenging to evaluate its effectiveness comprehensively. 2) For the synthetic data to real-scanned data (S2R) task, many analogous classes have been removed without clear justification. 3) Relying solely on accuracy as a metric means that incremental classes may not receive the emphasis they warrant. In light of these observations, we introduce a new benchmark: FSCIL3D-XL. Details of this benchmark will be discussed in the subsequent sections.
@@ 4.1 Task Setting
S2R Task. Chowdhury et al. [3]’s benchmark unjustifiably excludes certain classes that, while similar, are distinct in fact. For instance, semantically similar classes like “handbag” and “bag”, or shape-analogous classes like “ashcan” and “can”, are filtered out. By contrast, we retain all available classes in our benchmark. Recognizing the subtle differences between such similar classes is undoubtedly challenging, but we believe it is essential for a comprehensive evaluation. We continue to use ShapeNet as the base dataset and have selected CO3D [23] as the incremental dataset. From CO3D, we exclude 9 overlapping classes, resulting in 41 distinct classes designated for the incremental tasks.
For a more detail of the task settings in FSCIL3D-XL, please refer to Sec. 3 of Suppl..
@@ 4.2 Evaluation metrics
3D datasets present challenges distinct from their 2D counterparts. Firstly, there is a pronounced imbalance in the scale of different classes within 3D datasets. For instance, in ShapeNet, the “chair” class boasts over 1,000 training samples, whereas “birdhouse” has a mere 15. Secondly, incremental samples in 3D datasets, particularly those sourced from real-scanned data, tend to be more intricate. These novel classes, given their complexity, demand heightened attention in the FSCIL task. Relying solely on the accuracy metric, denoted as ${\rm Acc}$, would inadequately addresses these challenges. As a result, while we retain the 2D evaluation metrics ${\rm[Acc_{i}]_{B}}$ (accuracy of each session) and ${\Delta}$ [27] (relative accuracy dropping rate, where ${\rm\Delta=\frac{|Acc_{B}-Acc_{1}|}{Acc_{1}}}$), we also introduce new evaluation metrics specifically designed to address the aforementioned issues.
@@
where $K_{b}=\sum_{i=1}^{b}|Y_{b}|$ donates the number of visible classes for the $b$-th task, and $Acc_{i}$ donates the accuracy of the $i$-th class.
@@
where $P_{b}$ donates the predicted labels for test stage of $b$-th task and $B$ donates the number of task. F-Score. F-Score for FSCIL task (${\rm F_{FSCIL}}$, ${\rm F}$ for short): The network needs to be plastic to learn new knowledge from the current task, and it also needs to be stable to maintain knowledge learned from previous tasks. To ensure that the model is not overly biased towards either of these aspects, we refer to ${\rm F_{score}}$ and propose ${\rm F_{FSCIL}}$, which balances the plasticity and stability of the evaluation network. It can be calculated using the following formula:
@@
Finally, the 3D dataset is still being refined, so a large number of new datasets will likely continue to be produced. We take this into account and design our benchmark to be more flexible and modular in adding datasets, i.e., the desired FSCIL dataset can be obtained by simply transmitting the parameters and datasets to our generator. Additionally, The benchmark will be open-resource.
@@ 5 Experiments
In this section, we first provide more detailed information about our models and experiments. Then, we present and analyze the results of the comparison experiments. Finally, we present our ablation experiment to verify the components of our model.
@@
Table 1: Quantitative results on the S2S task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
@@
SimpleCIL-3D (ours) - 90.4 88.0 85.6 85.0 84.1 78.4 80.1 68.2 11.4 73.7
@@
$Joint$ FILP-3D - 90.6 89.1 88.5 87.8 87.4 87.6 86.8 79.9 4.2 83.2
80.0 78.5 79.0 77.2 77.5 77.7 77.3 79.9 3.4 78.6
Table 2: Quantitative results on the S2R task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
@@
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
$Joint$ FILP-3D 90.0 89.2 89.0 88.4 88.1 87.7 87.2 86.9 85.9 84.6 83.1 83.1 54.9 7.7 66.1
79.4 77.2 76.3 75.9 75.5 73.5 73.3 73.4 70.2 69.3 68.2 68.2 60.0 14.1 63.8
5.1 Implementation Details
We choose CLIP’s ViT-B/32 [21] as our pre-trained model, replace its visual encoder with CLIP2Point’s pre-trained depth encoder [11] and adopt CLIP2Point’s proposed rendering approach. As for the 3D encoder, we use DGCNN [34] pre-trained on ShapeNet, which follows OcCo [32]. Dimensions of both image features and text features are 512. We use SVD [9] as our dimensionality reduction method. We extract principal components (242 out of 512) based on the base task, which retains 95% energy of the principal components. We set the temperature parameter of infoNCE $\tau$ and balance parameter $\alpha$ in Eq. 8 to 0.1 and 1.0 respectively. For training, we use ADAM weight decay optimizer [13]. We set the learning rate to $1\times 10^{-3}$ and the weight decay to $1\times 10^{-4}$. Training in the base task and incremental take 10 epochs and 20 epochs respectively. For incremental tasks, we randomly select 5 samples/classes for training and allow 1 exemplar/class from previous tasks to be used as memory. The batch size is 32.
5.2 Experimental Results
We conduct extensive experiments on FSCIL3D-XL’s S2S task (Table 1), FSCIL3D-XL’s S2R task (Table 2), and Chowdhury et al. [3]’s benchmark (Sec. 6 of Suppl.). To make a comprehensive comparison, we choose the following models: 1) $Joint$: Models are trained with samples from all currently visible classes, which can represent their upper bound in FSCIL tasks. 2) One 3D SOTA approach developed by Chowdhury et al. [3]. 3) Two 2D SOTA approaches FACT [46] and BiDist [45]. FACT uses manifold-mixup [31] and virtual prototypes to ensure forward compatibility. BiDist uses knowledge distillation to retain knowledge. We use official codes to reproduce these methods, keeping all the original parameters unchanged. The only difference is that we replace the CNN-based network with CLIP2Point’s depth encoder to allow 3D application. For visualization results, more comparison results (including ULIP [37] and PointCLIP [43]) and effect of our proposed metrics, please refer to Sec. 4,7,8 of Suppl.
Experiments on the S2S task of FSCIL3D-XL. Table 1 reports the results of the synthetic data to the synthetic data task in our benchmark. BiDist [45] additionally employs a weighted sum of distillation loss for optimization, along with a relatively higher loss weight assigned to it. Therefore, BiDist tends to preserve existing knowledge and sacrifice the learning capacity for new classes, leading to a relatively low $\rm NCAcc$. Moreover, the S2S task has a limited proportion of incremental data, BiDist can achieve top $\rm Acc$ in over half of the tasks by keeping old knowledge. Contrary to it, FILP-3D has a trade-off between the base classes and the novel classes. The results also show that our metrics are reasonable for the evaluation of 3D FSCIL.
Experiments on the S2R task of FSCIL3D-XL. Table 2 reports the results of the synthetic data to the real-scanned data task in our benchmark. Considering the large domain gap between synthetic data and real-scanned data, it is even harder to acquire knowledge from new classes based on previous knowledge of synthetic shapes. We notice that BiDist witnesses a catastrophic performance decline from task 1 to task 2, which further demonstrates that BiDist tends to retain old knowledge and sacrifice the acquisition of new knowledge. In contrast, SimpleCIL-3D has already outperformed the other three methods, which proves that prior knowledge from pre-training models is effective for few-shot incremental tasks. However, we notice that its $\rm NCAcc$ is much lower than ours, indicating that SimpleCIL-3D performs poorly in real-scanned data. With the Spatial Noise Compensator, our FILP-3D achieves a state-of-the-art performance, almost close to its upper bound.
Table 3: Ablation studies on the S2R task of FSCIL3D-XL. For no contrastive learning (CL) case, training loss $\mathcal{L}^{b}=\mathcal{L}^{b}_{cls}$
. RFE SNC CL 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$ 86.7 71.2 49.6 17.9 58.5 78.9 56.3 49.0 28.6 52.4 90.4 75.6 50.8 16.3 60.8 80.0 58.3 51.0 27.1 54.4 90.1 73.6 54.5 18.3 62.6 79.3 56.7 56.1 28.5 56.4 89.8 74.0 57.8 17.6 64.9 78.7 56.7 58.0 28.0 57.3 90.0 74.6 60.0 17.1 66.9 79.4 57.3 59.9 27.8 58.6
5.3 Ablation Studies
To evaluate the effectiveness of our proposed modules and components, we conduct several ablation experiments. We verify the RCS, the Redundant Feature Eliminator on the S2S task (Sec. 2 of Suppl.), and further verify the Redundant Feature Eliminator, the Spatial Noise Compensator and contrastive learning on the S2R task (Table 3).
Effect of Redundant Feature Eliminator. For no RFE case (line 1 and 3), we use cosine similarity between $\mathbf{f}^{g}$ and ${\rm\mathbf{F}}^{t}$ to calculate logits. One can notice that the model with RFE (line 2) can only slightly outperform SimpleCIL-3D (line 1) in $\rm NCAcc$. We analyze the reason as follows: The huge domain gap leads to some semantic features available in synthetic data can not be extracted efficiently on real-scanned data. Adding some real-scanned data samples in the principal component extraction stage as prior knowledge may effectively alleviate this symptom. Ablation study for RFE in the S2S task (Sec. 2 of Suppl.) can demonstrate a more reasonable comparison and prove the effectiveness of our method in solving redundant feature problems.
Effect of Spatial Noise Compensator. Let’s compare the results of the line 2 with the line 4. On the one hand, there is a slight decrease in $\rm Acc$, which indicates that the 3D module is not as capable as the multi-view model with 2D prior knowledge. It is suitable only for supplementing the information in the 3D FSCIL task. On the other hand, there is a significant increase in the incremental part of the evaluation metric, indicating that the 3D module can complement the global information that the multi-view model lacks and enhance the anti-interference ability.
6 Conclusion
In this paper, we proposed FILP-3D, a 3D few-shot class-incremental learning framework with pre-trained V-L models. Specifically, we introduced a V-L pre-trained model CLIP to the 3D FSCIL task. To guarantee that CLIP performs well in the 3D FSCIL task, we proposed a Redundant Feature Eliminator to eliminate redundant features without stretching semantic information and a Spatial Noise Compensator to complement noise information. To comprehensively evaluate the 3D FSCIL task, we further proposed a new benchmark FSCIL3D-XL, which retains all classes and introduces reasonable evaluation metrics. Extensive experiments demonstrate that our FILP-3D achieves state-of-the-art performance in all available benchmarks. Results from ablation studies further verify the effectiveness of the proposed components.
Supplementary Material
In the supplementary materials, we introduce a detailed discussion of RCS in Section 7 and conduct ablation experiments to demonstrate the effectiveness of RCS and RFE in Section 8 . We provide more details about the task settings of our proposed benchmark FSCIL-3D XL in Section 9 as a supplement to our proposed benchmark. We provide visualizations in Section 10 for a more intuitive comparison between our model and other models (or ablation results). We elaborate on the preprocessing details of the principal component $\mathbf{V}$ in Section 11 to demonstrate the details of our model. We present experiments conducted on Chowdhury et al.’s benchmark in Section 12 and comparison experiments with more methods in Section 13 to demonstrate the superiority of our model as well as the generality and effectiveness of our proposed modules. Finally, by analyzing the experimental results, we demonstrate how our proposed metrics offer a more comprehensive evaluation of 3D FSCIL models in Section 14.
7 Detail Discussion for RCS
$\mathbf{V}$ is the principal components extracted from the base task, $\mathbf{v}_{i}$ is the $i$-th normalized principal component. $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$ are the projection of $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ onto the $\mathbf{V}$ vector space, formally expressed as the following equation:
$\mathbf{f}^{d}=\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}+\mathbf{R}^{d}$ (13)
$\mathbf{F}^{t}_{k}=\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k}$ (14)
where $M$ is the number of principal components, $\widetilde{\mathbf{f}}^{d}_{i}$ and $\mathbf{F}^{t}_{k,i}$ are the $i$-th dimension of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, respectively, and $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$ are the components of $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ which are not in the $\mathbf{V}$ vector space. $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$ are low-variance components (due to the nature of principal component analysis) and thus are redundant components according to discussions in Section 3.4.1.
Each $\mathbf{v}_{i}$ is an orthogonal vector and orthogonal to vectors outside the $\mathbf{V}$ vector space, e.g., $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$. Therefore $\forall i,j,k,\ i\neq j$ The following equation holds:
$\mathbf{v}_{i}\mathbf{v}^{T}_{j}=0,\ \mathbf{v}_{i}\mathbf{v}^{T}_{i}=1$ (15)
$\mathbf{R}^{d}\mathbf{v}_{i}^{T}=0,\ \mathbf{v}_{i}\mathbf{R}^{t^{T}}_{k}=0$ (16)
Then, we can derive the following equation:
$l_{k}$ $=\frac{\mathbf{f}^{d}\mathbf{F}^{t^{\mathrm{T}}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{(\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}+\mathbf{R}^{d})(\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k})^{T}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}\sum_{j=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,j}\mathbf{v}_{j}^{T}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$+\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}\mathbf{R}^{t^{T}}_{k}+\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{R}^{d}\mathbf{v}_{i}^{T}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\widetilde{\mathbf{F}}^{t}_{k,i}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{T}}_{k}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
Now we can observe that there are two terms in the numerator of this equation. $\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{T}}_{k}$ represents semantic information and should be retained, and $\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}$ is the multiplication of redundant feature vectors that should be compressed to keep only semantic information. Thus, the logit of the class $k$ can be calculated in a modified way as follows to eliminate redundant features:
$\widetilde{l}_{k}=\frac{\widetilde{\mathbf{f}}^{g}\widetilde{\mathbf{F}}^{t^{T}}_{k}}{\|\mathbf{f}^{g}\|\|\mathbf{F}^{t}_{k}\|}.$ (17)
8 Extra Ablation Studies
@@
Table 4: Ablation studies on the S2S task of FSCIL3D-XL

### v2

@@ 5 Benckmark for 3D FSCIL task
Studies on 3D FSCIL benchmarks are still in the early stages. The sole benchmark introduced by Chowdhury et al. [8] has several notable limitations: 1) In the synthetic data to synthetic data (S2S) task set forth by Chowdhury et al. [8], the total number of classes is limited. This restricts the capacity of the model for incremental learning and makes it challenging to evaluate its effectiveness comprehensively. 2) For the synthetic data to real-scanned data (S2R) task, many analogous classes have been removed without clear justification. 3) Relying solely on accuracy as a metric means that incremental classes may not receive the emphasis they warrant. In light of these observations, we introduce a new benchmark: FSCIL3D-XL. Details of this benchmark will be discussed in the subsequent sections.
@@ 5.1 Task Setting
S2R Task. Chowdhury et al. [8]’s benchmark unjustifiably excludes certain classes that, while similar, are distinct in fact. For instance, semantically similar classes like “handbag” and “bag”, or shape-analogous classes like “ashcan” and “can”, are filtered out. By contrast, we retain all available classes in our benchmark. Recognizing the subtle differences between such similar classes is undoubtedly challenging, but we believe it is essential for a comprehensive evaluation. We continue to use ShapeNet as the base dataset and select CO3D [13] as the incremental dataset. From CO3D, we exclude 9 overlapping classes, resulting in 41 distinct classes designated for the incremental tasks.
@@ 5.2 Evaluation metrics
3D datasets present challenges distinct from their 2D counterparts. Firstly, there is a pronounced imbalance in the scale of different classes within 3D datasets. For instance, in ShapeNet, the “chair” class boasts over 1,000 training samples, whereas “birdhouse” has a mere 15. Secondly, incremental samples in 3D datasets, particularly those sourced from real-scanned data, tend to be more intricate. These novel classes, given their complexity, demand heightened attention in the FSCIL task. Relying solely on the micro accuracy metric, denoted as ${\rm Acc}$, would inadequately address these challenges. As a result, while we retain the 2D evaluation metrics ${\rm Acc_{i}}$ (micro accuracy of $i$-th session, where ${\rm Acc_{i}}=\rm{\frac{CP}{TP}}$. $\rm{CP}$ and $\rm{TP}$ denote the correct/total predictions in $i$-th session respectively.) and ${\Delta}$ [35] (relative accuracy dropping rate, where ${\rm\Delta=\frac{|Acc_{B}-Acc_{1}|}{Acc_{1}}}$), we also introduce new evaluation metrics specifically designed to address the aforementioned issues.
@@
where $K_{b}=\sum_{i=1}^{b}|Y_{b}|$ denotes the number of visible classes for the $b$-th task, and $Acc_{i}$ denotes the accuracy of the $i$-th class.
@@
where $P_{b}$ denotes the predicted labels for test stage of $b$-th task and $B$ denotes the number of task.
F-Score. F-Score for FSCIL task (${\rm F_{FSCIL}}$, ${\rm F}$ for short): The network needs to be plastic to learn new knowledge from the current task, and it also needs to be stable to maintain knowledge learned from previous tasks. To ensure that the model is not overly biased towards either of these aspects, we refer to ${\rm F_{score}}$ and propose ${\rm F_{FSCIL}}$, which balances the plasticity and stability of the evaluation network. It can be calculated using the following formula:
@@
Finally, the 3D dataset is still being refined, so a large number of new datasets will likely continue to be produced. We take this into account and design our benchmark to be more flexible and modular in adding datasets, i.e., the desired FSCIL dataset can be obtained by simply transmitting the parameters and datasets to our generator. Additionally, The benchmark has been open-resource.
@@ 6 Experiments
In this section, we begin by furnishing a more comprehensive exposition of our models and experimental setup. Subsequently, we proceed to present and analyze the results of the comparative experiments. Finally, we showcase ablation experiments and visualizations to substantiate the effectiveness of our proposed modules. Please note that we only utilize SNC during incremental learning with real-scanned data, such as in the S2R task. In cases where SNC is not applied such as the S2S task, we employ cosine similarity between $\mathbf{f}^{2D}$ and ${\rm\mathbf{F}}^{t}$ for classification.
Table 2: The number of samples $n_{b}$ in the training and testing sets for the S2S and S2R tasks
Task Type The number of samples $n_{b}$
0 1 2 3 4 5 6 7 8 9 10 11
ShapeNet2CO3D. [8] train 26287 64 69 74 79 84 89 94 99 104 109 -
test 6604 6793 6959 7205 7333 7492 7646 7816 7957 8117 8336 -
S2S train 41943 75 79 83 87 91 95 - - - - -
test 10517 10663 10889 11195 11435 11595 11835 - - - - -
S2R train 41943 75 79 83 87 91 95 99 103 107 111 100
test 10517 10867 11020 11395 11705 11897 12226 12505 12718 13022 13313 13404
6.1 Implementation Details
We select CLIP’s ViT-B/32 [12] as our pre-trained model, replacing its visual encoder with CLIP2Point’s pre-trained depth encoder [17] and adopting CLIP2Point’s proposed rendering approach. The number of views $N$ is set to 6. For the 3D encoder, we utilize DGCNN [21] pretrained on ShapeNet, following the methodology of OcCo [36]. Both image and text embedding dimensions $C$ are 512. We employ SVD [37] as our dimensionality reduction technique, extracting 242 principal components based on the base task (i.e. $M$ = 242), which retains 95% energy of the principal components. The temperature parameter of infoNCE and $\mathcal{L}_{align}$ are set to 0.7 and 1.0 respectively. The balance parameter $\alpha$ and $\beta$ in $\mathcal{L}_{align}$ are set to 0.25. For training, we utilize the ADAM weight decay optimizer [38], with the learning rate set to $1\times 10^{-3}$ and the weight decay to $1\times 10^{-4}$. Training in the base task and incremental tasks lasts for 10 and 20 epochs, respectively. In incremental tasks, we randomly select 5 samples/classes for training and allow 1 exemplar/class from previous tasks to be utilized as memory. The number of samples $n_{b}$ in the training and testing sets for the S2S and S2R tasks is shown in Table  2. The batch size is set to 32. All of our experiments are conducted on 1 A6000 GPU using PyTorch.
6.2 Computational Cost
An iteration (it) processes one batch, which consists of 32 samples. During the training of base classes, the GPU memory consumption is 7814 MiB, with an average inference speed of 1.34 it/s. During incremental training, the GPU memory consumption increases to 12838 MiB, with an average inference speed of 3.41 s/it. During test inference, the GPU memory consumption remains at 7776 MiB, with an average inference speed of 2.65 it/s.
6.3 Experimental Results
Compared methods. To make a comprehensive comparison, we choose the following models: 1) $Joint$: Models are trained with samples from all currently visible classes, which can represent their upper bound in FSCIL tasks. 2) One 3D FSCIL SOTA approach developed by Chowdhury et al. [8]. 3) Two 3D few-shot SOTA approaches PointCLIP [15] and ULIP [3] which also utilize CLIP as backbone. We compare our model with these two models to validate the effectiveness of our model’s design in addressing continuous-learning-related issues. We further introduce PointCLIP++ (PointCLIP backbone + RFE + SNC) to validate the generalization of our proposed modules. 4) Two 2D SOTA approaches FACT [5] and BiDist [7]. FACT uses manifold-mixup [39] and virtual prototypes to ensure forward compatibility. BiDist uses knowledge distillation to retain knowledge. We use official codes to reproduce these methods, keeping all the original parameters unchanged. The only difference is that we replace the CNN-based network with CLIP2Point’s depth encoder to allow 3D application.
Notice that joint training suffers from the long-tail issue on new classes but can be properly solved on the FSCIL framework. Therefore, joint training results are not exactly the upper limit of $\rm NCAcc$ and $\rm F$. Therefore we exclude $\rm NCAcc,F$ metrics of $joint$ in the experiment table.
@@ 6.3 Experimental Results
Table 3: Quantitative results on Chowdhury et al. [8]’s ShapeNet2CO3D benchmark. Bold denotes the best performance, $joint$ serves solely as an upper reference limit in Chowdhury et al. [8]’s work and is not involved in the comparison.
Method Acc. in each session $\uparrow$ ${\rm\Delta}\downarrow$
0 1 2 3 4 5 6 7 8 9 10
FACT [5] 81.4 76.0 70.3 68.1 65.8 63.5 63.0 60.1 58.2 57.5 55.9 31.3
baseline (ours) 85.4 81.3 80.1 77.9 75.1 73.8 72.3 70.6 69.3 67.1 65.6 23.2
FILP-3D (ours) 87.3 81.5 82.3 80.1 78.5 76.4 76.8 74.2 71.6 71.7 70.1 19.7
Experiments on Chowdhury et al. [8]’s benchmark. We report the performance on Chowdhury et al. [8]’s ShapeNet2CO3D benchmark in Table 3. Chowdhury et al. [8]’s work does not provide any detailed data partitions other than ShapeNet2CO3D. Consequently, we are unable to follow them and conduct further experiments and comparisons, such as incremental within a single dataset. As shown in Table 3, FILP-3D outperforms the other two methods. After the final task, our accuracy only drops to 70.1%, which surpasses the upper bound of Chowdhury et al. [8]’s method. This demonstrates the vast potential of CLIP in the 3D FSCIL task. However, we note that the results in this benchmark cannot distinctly indicate whether the performance degradation of  Chowdhury et al. [8] is caused by the new class or the old classes. We cannot conclude more observations based on it.
Table 4: Quantitative results on the S2S task. For each set of results, the micro/macro average are presented at the top/bottom respectively. Bold denotes the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
@@
ULIP [3] CVPR’23 86.3 83.3 80.3 75.8 72.8 62.9 65.1 78.3 24.6 71.1
85.4 80.5 76.3 71.4 69.3 67.6 64.9 77.2 24.0 70.5
PointCLIP [15] CVPR’22 88.3 85.8 81.0 73.9 72.0 71.9 67.7 54.3 23.3 60.3
75.8 71.8 67.5 63.6 62.1 59.9 56.9 54.9 24.9 55.9
PointCLIP++ - 88.8 86.3 82.3 80.0 77.5 74.0 70.3 60.7 20.8 65.1
76.2 71.7 68.9 67.5 64.2 61.1 59.8 61.6 21.5 60.7
baseline (ours) - 90.4 88.0 85.6 85.0 84.1 78.4 80.1 68.2 11.4 73.7
@@
$Joint$ FILP-3D - 90.6 89.1 88.5 87.8 87.4 87.6 86.8 - 4.2 -
80.0 78.5 79.0 77.2 77.5 77.7 77.3 - 3.4 -
Experiments on the S2S task of FSCIL3D-XL. Table 4 reports the results of the synthetic data to the synthetic data task in our benchmark, whilch mainly reflect models’ capability to mitigate issues like overfitting or catastrophic forgetting.
FACT [5], Chowdhury et al. [8]’s work, and ULIP lack the shape-relative prior knowledge and therefore perform worse than methods that use CLIP as the backbone, such as PointCLIP [15] and baseline. This demonstrates that V-L PTMs are effective in addressing the dual challenges of forgetting and overfitting, not only in 2D FSCIL tasks but also in 3D FSCIL tasks.
BiDist [7] additionally employs a weighted sum of distillation loss for optimization, along with a relatively higher loss weight assigned to it. Therefore, BiDist tends to preserve existing knowledge and sacrifices the learning capacity for new classes, leading to the highest $\rm Acc$ and lowest $\rm NCAcc$. Contrary to it, FILP-3D has a trade-off between the base classes and the novel classes. The results also show that our metrics are reasonable for the evaluation of 3D FSCIL.
The PointCLIP [15] architecture is similar to the baseline, hence they exhibit comparable performance. PointCLIP++ and FILP-3D show significant performance improvements over PointCLIP and the baseline respectively, particularly for new classes ($\rm{NCAcc}$). This phenomenon highlights the importance of the feature space alignment issue addressed by RFE, which plays a crucial role in enabling V-L PTMs to fully leverage their potential in 3D FSCIL tasks.
Table 5: Quantitative results on the S2R task. For each set of results, the micro/macro average are presented at the top/bottom respectively. Bold denotes the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
@@
FILP-3D (ours) 90.2 87.1 85.9 84.4 83.4 82.3 80.4 79.3 78.1 76.7 74.6 74.4 60.4 17.5 66.7
79.5 75.3 74.8 71.6 69.6 67.9 65.3 62.7 61.7 60.0 58.2 57.7 61.8 27.4 59.7
$Joint$ FILP-3D 90.0 89.2 89.0 88.4 88.1 87.7 87.2 86.9 85.9 84.6 83.1 83.1 - 7.7 -
79.4 77.2 76.3 75.9 75.5 73.5 73.3 73.4 70.2 69.3 68.2 68.2 - 14.1 -
Experiments on the S2R task of FSCIL3D-XL. Table 5 reports the results of the synthetic data to the real-scanned data task in our benchmark. Considering the large domain gap between synthetic data and real-scanned data, it is even harder to acquire knowledge from new classes based on previous knowledge of synthetic shapes.
By observing the experiment results presented in Table 5, we notice a catastrophic decline in FACT [5]’s performance in terms of $\rm MAcc$ when learning real-scanned data. This may be because The feature space trained by the manifold-mixup [39] method is relatively fragile in the aspect of structure, thus it is easy to overfit the noise information during the incremental process, especially for real-scanned data. Solely relying on metrics proposed by Chowdhury et al., we are unable to identify the aforementioned issues and then perform the analysis.
We also notice that, unlike the experiment on the S2S task, BiDist [7] witnesses a catastrophic performance decline from task 1 to task 2, which further demonstrates that BiDist is unable to balance the retaining of old knowledge and the acquisition of new knowledge in scenarios with a large domain gap. In contrast, our baseline has already outperformed the other three methods, which proves that prior knowledge from pre-training models is effective for few-shot incremental tasks.
Comparing ULIP with our model, it’s evident that ULIP exhibits outstanding learning capabilities in novel classes. However, the abundance of trainable parameters and severely inadequate data make it extremely challenging to avoid overfitting during the FSCIL process, ultimately resulting in a high dropping rate $\rm\Delta$ and limited performance. Consequently, we conclude that a frozen 2D PTM as the backbone is more suitable for 3D FSCIL.
The PointCLIP also exhibits comparable performance to the baseline because of similar architecture. Experimental results demonstrate that PointCLIP++ outperforms PointCLIP significantly across various metrics, especially in $\rm NCAcc$. This validates the generality and effectiveness of our proposed modules.
For our FILP-3D, the 2D PTM and simple adapters provide FILP-3D with good generality and adaptability. On the other hand, RFE and SNC focus on addressing specific issues related to applying multi-view 3D models to FSCIL tasks, thereby achieving SOTA performance.
Effect of our proposed metrics. By observing the experiment results presented in Table 4, a notable deterioration in Chowdhury et al.’s [8] performance is observed concerning $\rm MAcc$ during incremental learning, accompanied by subpar results in the $\rm NCAcc$ metric. This indicates their ineffectiveness in acquiring new knowledge efficiently. Solely relying on micro Acc. and dropping rate $\Delta$, we are unable to identify the aforementioned issues. Similarly, without $\rm NCAcc$, Bidist’s inability to effectively learn new classes also can not be revealed. Similarly, without the $\rm NCAcc$ metric in Table 5, Chowdhury et al.’s terrible performance in learning real scanned new classes also can not be noticed.
Table 6: Compare RFE with other methods.
Method 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
APE 90.3 72.8 53.8 19.4 61.9
79.3 56.5 55.7 28.8 56.1
DomainDrop 90.3 72.7 54.4 19.5 62.2
79.9 55.9 56.1 30.0 56.0
FILP-3D 90.2 74.4 60.4 17.5 66.7
79.5 57.7 61.8 27.4 59.6
Comparsion between RFE with other channel mask methods. We conduct comparison experiments between RFE with APE [9] and DomainDrop [40] in the S2R task to reveal the superiority of RFE. APE utilizes textual features of classes to filter out redundant channels. However, the number of classes in the base task (55) is limited compared to APE (1000), making it prone to mistakenly masking semantic features. The random drop of domain-sensitive channels proposed by DomainDrop with a 33% drop rate is too aggressive and unstable for the few-shot and incremental setting, leading to overfitting and forgetting. These methods are primarily tailored for domain adaptation tasks and consequently exhibit limited performance due to their failure to address the challenges inherent in FSCIL. Contrary to them, our proposed RFE fully leverages abundant data from base classes to preprocess more general principal components, combined with the RCS to avoid DR compromising semantic integrity. According to Table 6, our RFE outperforms APE and DomainDrop in terms of almost all metrics.
@@
Table 7: Ablation studies on the S2S task of FSCIL3D-XL. The ablation experiments for RFE and RCS were conducted in this set of experiments.